Imagine that you are the supply chain manager for the Magic Widget company and you need to measure your supply chain performance. The chart shows the financial variables that you will need to perform your task.

Financial Variables
Financial Variables Amount in $ Billions
Inventory Investment 10.8
Total Assets 38.5
Cost of Goods Sold 13.1
Raw Material Inventory .85
Work-in-progress Inventory .15
Finished Goods Inventory .85

Compute the percentage of assets committed to inventory and inventory turnover. Round your answers to the first decimal place.

The total assets committed to inventory is__________% and the inventory turnover for the firm is _________ times.
Answer #1

The total assets committed to inventory = (Inventory invested / Total assets) * 100

= 10.8 / 38.5 * 100 = 28%

Turnover = Cost of goods sold / Inventory

= 13.1 / 10.8 = 1.21 times

Do give a positive rating if you think this answer helped